Violet quit her job on Bay street earning $91,000 a year to start her own financial planning business. She earned $171,000 in revenue. She paid herself $100,000 and spent $30,000 on materials, rent, client retention etc. Based on this information, what were her economic profit, economic costs and her accounting costs?
Answer #1

Opportunity cost (Economic) = $ 91,000

Revenue = $ 171,000

Accounting cost = $ 30,000

Accounting profit = Revenue - Explicit cost = 171,000-30000 = $141,000

Salary to owner = $ 100,000(Implicit cost)

Economic profit =Accounting profit - implicit cost - opportunity cost

Economic profit = 141,000 - 100,000 - 91,000 = - $ 50,000

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me please help. Thank you.